# Gorenstein singularities with $\mathbb{G}_{m}$-action and moduli spaces of holomorphic differentials

### v1

@@ 5.2. The stratum $\mathbb{P}\mathcal{H}(2,2)^{\operatorname{odd}}$
[…]
@@
Since $x$, $y$, and $z$ have weights $2$, $2$, and $3$ under the $\mathbb{G}_{m}$-action, respectively, the standard projectivization of the corresponding Gorenstein curve $Y$ as well as its deformations $(C,p_{1},p_{2})\in\mathbb{P}\mathcal{H}(2,2)^{\operatorname{odd}}$ can be embedded in the weighted projective space $\mathbb{P}(2,2,3,1)$, where the divisor at infinity is cut out by the hyperplane of weight $1$ as $p_{1}+p_{2}$.
@@
The nontrivial weights are $1$, $1$, and $3$. Combining with Proposition 4.1 (iv) and (11), we thus obtain that
@@ 5.3. The stratum $\mathbb{P}\mathcal{H}(2,1,1)$
[…]
We want to describe the generators of $\mathfrak{m}/\mathfrak{m}^{2}\pmod{\widetilde{\mathfrak{m}}^{4}}$. Up to rescaling the parameters, we can assume that the linear generator is $0\oplus t_{2}\oplus t_{3}\in R$. Since $0\oplus t_{2}^{2}\oplus 0\not\in R$ and $0\oplus 0\oplus t_{3}^{2}\not\in R$, the other quadratic generators can be written as $t_{1}^{2}\oplus 0\oplus 0\in R$ and $t_{1}^{3}\oplus t_{2}^{2}\oplus 0\in R$. It follows that $0\oplus t_{2}^{3}\oplus 0\in R$, $0\oplus 0\oplus t_{3}^{3}\in R$, and hence $\alpha_{3}=1$ is confirmed.
@@
$(xy,y(z^{2}-y^{3}),xz(z-x^{2}),z^{2}-y^{3}-x^{2}z).$
Since $x$, $y$, and $z$ have weights $3$, $4$, and $6$ under the $\mathbb{G}_{m}$-action, respectively, the standard projectivization of the corresponding Gorenstein curve $Y$ as well as its deformations $(C,p_{1},p_{2},p_{3})\in\mathbb{P}\mathcal{H}(2,1,1)$ can be embedded in the weighted projective space $\mathbb{P}(3,4,6,1)$, where the divisor at infinity is cut out by the hyperplane of weight $1$ as $\frac{1}{2}p_{1}+\frac{1}{3}p_{2}+\frac{1}{3}p_{3}$.
@@
The nontrivial weights are $2$, $3$, and $6$. We thus obtain that
@@ 5.4. The stratum $\mathbb{P}\mathcal{H}(5,1)$
[…]
@@
$(xz-y^{2},x^{2}y-z^{2},y^{5}-z^{4}).$
Since $x$, $y$, and $z$ have weights $3$, $4$, and $5$ under the $\mathbb{G}_{m}$-action, respectively, the standard projectivization of the corresponding Gorenstein curve $Y$ as well as its deformations $(C,p_{1},p_{2})\in\mathbb{P}\mathcal{H}(5,1)$ can be embedded in the weighted projective space $\mathbb{P}(3,4,5,1)$, where the divisor at infinity is cut out by the hyperplane of weight $1$ as $p_{1}+\frac{1}{3}p_{2}$.
@@
The nontrivial weights are $1$, $2$, $3$, and $6$. We thus obtain that
@@
[…]
We want to describe the generators of $\mathfrak{m}/\mathfrak{m}^{2}\pmod{\widetilde{\mathfrak{m}}^{6}}$. Since the exponents of nonzero terms of $t_{1}$ and $t_{2}$ in every (non-decomposable) generator are proportional as $5:3$, we have $\alpha_{1}=2$, and hence $\alpha_{2}=1$ or 00.
@@
    Suppose $\alpha_{2}=0$. Then $\alpha_{4}=0$ and $\alpha_{3}=1$. In this case, the gap sequence is [2 0 1 0 1]delimited-[]20101[2\ 0\ 1\ 0\ 1][ 2 0 1 0 1 ]. It follows that $t_{1}^{2}\oplus 0\in R$, $0\oplus t_{2}^{2}\in R$, and $at_{1}^{5}\oplus t_{2}^{3}\in R$ for some $a\neq 0$. Up to rescaling $t_{1}$, we can assume that $a=1$ and $t_{1}^{5}\oplus t_{2}^{3}\in R$.
@@
    $(xy,x^{5}+y^{3}-z^{2}).$
@@
    Finally, we compute the weights and characters for this singularity. In this case, we have $\ell=15$, $a_{1}=3$, and $a_{2}=5$. Since $\omega_{C}^{\log}\sim 5p_{1}+3p_{2}$, it follows that
    $\mathcal{F}^{\lambda}H^{0}(C,\omega_{C}^{\log})=H^{0}(C,(5-\lceil\lambda/3\rceil)p_{1}+(3-\lceil\lambda/5\rceil)p_{2}).$
    Therefore, the filtration of $H^{0}(C,\omega_{C}^{\log})$ is given by
@@
    The nontrivial weights are $3$, $5$, $9$, and $15$. We thus obtain that
    $\chi_{1}^{\log}(4,2)^{\operatorname{even}}=32,\quad\chi_{2}^{\log}(4,2)^{\operatorname{even}}=152,\quad\alpha(4,2)^{\operatorname{even}}=\frac{14}{33}.$
@@
    Suppose $\alpha_{2}=1$. Then the gap sequence is [2 1 0 0 1]delimited-[]21001[2\ 1\ 0\ 0\ 1][ 2 1 0 0 1 ]. It follows that $t_{1}^{3}\oplus 0\in R$, $t_{1}^{5}\oplus t_{2}^{3}\in R$ (up to rescaling $t_{1}$), $t_{1}^{4}\oplus 0\in R$, and $0\oplus t_{2}^{4}\in R$. Since $t_{1}^{5}\oplus 0\not\in R$, the quadratic generator of $\mathfrak{m}/\mathfrak{m}^{2}$ must be $0\oplus t_{2}^{2}$.
@@
    $(xy,xw,yz-w^{2},zw-y^{3},x^{3}+y^{2}w-z^{2}).$
    Since $x$, $y$, $z$, and $w$ have weights $10$, $9$, $15$, and $12$ under the $\mathbb{G}_{m}$-action, respectively, the standard projectivization of the corresponding Gorenstein curve $Y$ as well as its deformations $(C,p_{1},p_{2})$ can be embedded in the weighted projective space $\mathbb{P}(10,9,15,12,1)$, where the divisor at infinity is cut out by the hyperplane of weight $1$ as $\frac{1}{3}p_{1}+\frac{1}{5}p_{2}$.
@@
    Finally, we compute the weights and characters for this singularity. In this case, we have $\ell=15$, $a_{1}=3$, and $a_{2}=5$. Since $\omega_{C}^{\log}\sim 5p_{1}+3p_{2}$, it follows that
    $\mathcal{F}^{\lambda}H^{0}(C,\omega_{C}^{\log})=H^{0}(C,(5-\lceil\lambda/3\rceil)p_{1}+(3-\lceil\lambda/5\rceil)p_{2}).$
    Therefore, the filtration of $H^{0}(C,\omega_{C}^{\log})$ is given by
@@
    The nontrivial weights are $3$, $5$, $6$, and $15$. We thus obtain that
    $\chi_{1}^{\log}(4,2)^{\operatorname{odd}}=29,\quad\chi_{2}^{\log}(4,2)^{\operatorname{odd}}=149,\quad\alpha(4,2)^{\operatorname{odd}}=\frac{79}{228}.$
@@ 5.6. The stratum $\mathbb{P}\mathcal{H}(3,3)^{\operatorname{nonhyp}}$
[…]
@@
Since $x$, $y$, and $z$ have weights $2$, $3$, and $3$ under the $\mathbb{G}_{m}$-action, respectively, the standard projectivization of the corresponding Gorenstein curve $Y$ as well as its deformations $(C,p_{1},p_{2})\in\mathbb{P}\mathcal{H}(3,3)^{\operatorname{nonhyp}}$ can be embedded in the weighted projective space $\mathbb{P}(2,3,3,1)$, where the divisor at infinity is cut out by the hyperplane of weight $1$ as $p_{1}+p_{2}$.
@@
The nontrivial weights are $1$, $1$, $2$, and $4$. We thus obtain that
@@ 5.7. The stratum $\mathbb{P}\mathcal{H}(3,2,1)$
[…]
@@
Since $x$, $y$, $z$, and $w$ have weights $6$, $8$, $9$, and $12$ under the $\mathbb{G}_{m}$-action, respectively, the standard projectivization of the corresponding Gorenstein curve $Y$ as well as its deformations $(C,p_{1},p_{2},p_{3})\in\mathbb{P}\mathcal{H}(3,2,1)$ can be embedded in the weighted projective space $\mathbb{P}(6,8,9,12,1)$, where the divisor at infinity is cut out by the hyperplane of weight $1$ as $\frac{1}{3}p_{1}+\frac{1}{4}p_{2}+\frac{1}{6}p_{3}$.
@@
The nontrivial weights are $3$, $4$, $6$, and $12$. We thus obtain that
@@ 5.8. The stratum $\mathbb{P}\mathcal{H}(2,2,2)^{\operatorname{odd}}$
[…]
@@
Since $x$, $y$, $z$, $w$, and $u$ have weights $2$, $2$, $2$, $3$, and $3$ under the $\mathbb{G}_{m}$-action, respectively, the standard projectivization of the corresponding Gorenstein curve $Y$ as well as its deformations $(C,p_{1},p_{2},p_{3})\in\mathbb{P}\mathcal{H}(2,2,2)^{\operatorname{odd}}$ can be embedded in the weighted projective space $\mathbb{P}(2,2,2,3,3,1)$, where the divisor at infinity is cut out by the hyperplane of weight $1$ as $p_{1}+p_{2}+p_{3}$.
@@
The nontrivial weights are $1$, $1$, $1$, and $3$. We thus obtain that

### v2

@@ 5.2. The stratum $\mathbb{P}\mathcal{H}(2,2)^{\operatorname{odd}}$
In this case, we have $g=3$, $b=2$, and $\delta=4$. The gap sequence is determined by $\alpha_{1}\ \alpha_{2}\ 1]$, where $\alpha_{3}=1$ is the last nonzero entry and $\sum_{i\geq 1}\alpha_{i}=3$. Let $y_{i}$ be the point at infinity in the rational branch $Y_{i}$ for $i=1,2$. The spin parity of this case is determined by $\dim H^{0}(Y,y_{1}+y_{2})\pmod{2}$, where $t_{1}^{n_{1}}\oplus t_{2}^{n_{2}}\in H^{0}(Y,y_{1}+y_{2})$ if and only if $n_{1}\leq 1$ and $n_{2}\leq 1$. Since we want $H^{0}(Y,y_{1}+y_{2})$ to be one-dimensional spanned by the constant function only, there is no linear generator in $R$. Consequently, $\alpha_{1}=2$ and the gap sequence is [2 0 1][2\ 0\ 1].
@@
This singularity is denoted by $T_{7}$ in Giusti’s classification of simple space curve singularities; see [Loo84a, (7.22)] and [Ste15, Table 2].
Since $x$, $y$, and $z$ have weights $2$, 22, and $3$ under the $\mathbb{G}_{m}$-action, respectively, the standard projectivization of the corresponding Gorenstein curve $Y$ as well as its deformations $(C,p_{1},p_{2})\in\mathbb{P}\mathcal{H}(2,2)^{\operatorname{odd}}$ can be embedded in the weighted projective space $\mathbb{P}(2,2,3,1)$, where the divisor at infinity is cut out by the hyperplane of weight $1$ as $p_{1}+p_{2}$.
@@
The nontrivial weights are $1$, 11, and $3$. Combining with Proposition 4.1 (iv) and (11), we thus obtain that
@@ 5.3. The stratum $\mathbb{P}\mathcal{H}(2,1,1)$
In this case, we have $g=3$, $b=3$, and $\delta=5$. The gap sequence is determined by $\alpha_{1}\ \alpha_{2}\ 1]$, where $\alpha_{3}=1$ is the last nonzero entry and $\sum_{i\geq 1}\alpha_{i}=3$. Note that the exponents of nonzero terms of $t_{1}$, $t_{2}$, and $t_{3}$ in every (non-decomposable) generator are proportional as $3:2:2$. Since each individual $t_{i}\not\in R$, there is only one linear generator in $R$, which implies that $\alpha_{1}=2$. Therefore, the gap sequence is [2 0 1][2\ 0\ 1].
We want to describe the generators of $\mathfrak{m}/\mathfrak{m}^{2}\pmod{\widetilde{\mathfrak{m}}^{4}}$. Up to rescaling the parameters, we can assume that the linear generator is $0\oplus t_{2}\oplus t_{3}\in R$. Since $0\oplus t_{2}^{2}\oplus 0\not\in R$ and $0\oplus 0\oplus t_{3}^{2}\not\in R$, the other quadratic generators can be written as $t_{1}^{2}\oplus 0\oplus 0\in R$ and $t_{1}^{3}\oplus t_{2}^{2}\oplus 0\in R$. It follows that $\oplus t_{2}^{3}\oplus 0\in R, 00$$\oplus 0\oplus t_{3}^{3}\in R$, and hence $\alpha_{3}=1$ is confirmed.
@@
$(xy,z^{2}-y^{3}-x^{2}z).$
This singularity is denoted by $T_{8}$ in Giusti’s classification of simple space curve singularities; see [Loo84a, (7.22)] and [Ste15, Table 2].
Since $x$, $y$, and $z$ have weights $3$, 44, and $6$ under the $\mathbb{G}_{m}$-action, respectively, the standard projectivization of the corresponding Gorenstein curve $Y$ as well as its deformations $(C,p_{1},p_{2},p_{3})\in\mathbb{P}\mathcal{H}(2,1,1)$ can be embedded in the weighted projective space $\mathbb{P}(3,4,6,1)$, where the divisor at infinity is cut out by the hyperplane of weight $1$ as $\frac{1}{2}p_{1}+\frac{1}{3}p_{2}+\frac{1}{3}p_{3}$.
@@
The nontrivial weights are $2$, 33, and $6$. We thus obtain that
@@ 5.4. The stratum $\mathbb{P}\mathcal{H}(5,1)$
In this case, we have $g=4$, $b=2$, and $\delta=5$. The gap sequence is determined by $\alpha_{1}\ \alpha_{2}\ \alpha_{3}\ \alpha_{4}\ \alpha_{5}\ 1]$, where $\alpha_{6}=1$ is the last nonzero entry. Moreover, $\sum_{i\geq 1}\alpha_{i}=4$ and $\alpha_{1},\alpha_{2},\alpha_{3}\neq 0$. Therefore, the only possibility is [1 1 1 0 0 1][1\ 1\ 1\ 0\ 0\ 1].
@@
$(xz-y^{2},x^{2}y-z^{2}).$
This singularity is denoted by $W_{9}$ in Giusti’s classification of simple space curve singularities; see [Loo84a, (7.22)] and [Ste15, Table 2].
Since $x$, $y$, and $z$ have weights $3$, 44, and $5$ under the $\mathbb{G}_{m}$-action, respectively, the standard projectivization of the corresponding Gorenstein curve $Y$ as well as its deformations $(C,p_{1},p_{2})\in\mathbb{P}\mathcal{H}(5,1)$ can be embedded in the weighted projective space $\mathbb{P}(3,4,5,1)$, where the divisor at infinity is cut out by the hyperplane of weight $1$ as $p_{1}+\frac{1}{3}p_{2}$.
@@
The nontrivial weights are $1$, 22, 33, and $6$. We thus obtain that
@@
In this case, we have $g=4$, $b=2$, and $\delta=5$. The gap sequence is determined by $\alpha_{1}\ \alpha_{2}\ \alpha_{3}\ \alpha_{4}\ 1]$, where $\alpha_{5}=1$ is the last nonzero entry and $\sum_{i\geq 1}\alpha_{i}=4$.
We want to describe the generators of $\mathfrak{m}/\mathfrak{m}^{2}\pmod{\widetilde{\mathfrak{m}}^{6}}$. Since the exponents of nonzero terms of $t_{1}$ and $t_{2}$ in every (non-decomposable) generator are proportional as $5:3$, we have $\alpha_{1}=2$, and hence $\alpha_{2}=1$ or 0.
@@
    Suppose $\alpha_{2}=0$. Then $\alpha_{4}=0$ and $\alpha_{3}=1$. In this case, the gap sequence is [2 0 1 0 1][2\ 0\ 1\ 0\ 1]. It follows that $t_{1}^{2}\oplus 0\in R$, $0\oplus t_{2}^{2}\in R$, and $at_{1}^{5}\oplus t_{2}^{3}\in R$ for some $a\neq 0$. Up to rescaling $t_{1}$, we can assume that $a=1$ and $t_{1}^{5}\oplus t_{2}^{3}\in R$.
@@
    In terms of these generators, the defining equations of this singularity are (xy, x^5 + y^3 - z^2). This singularity is denoted by $T_{9}$ in Giusti’s classification of simple space curve singularities; see [Loo84a, (7.22)] and [Ste15, Table 2].
@@
    Finally, we compute the weights and characters for this singularity. In this case, we have $\ell=15$, $a_{1}=3$, and $a_{2}=5$. Since $\omega_{C}^{\log}\sim 5p_{1}+3p_{2}$, it follows that F^λ H^0(C, ω_C^log) = H^0(C, (5 - ⌈λ/3⌉)p_1 + (3 - ⌈λ/ 5⌉) p_2). Therefore, the filtration of $H^{0}(C,\omega_{C}^{\log})$ is given by
@@
    The nontrivial weights are $3$, 55, 99, and $15$. We thus obtain that χ_1^log(4,2)^even = 32,  χ_2^log(4,2)^even = 152 ,  α(4,2)^even = 1433.
@@
    Suppose $\alpha_{2}=1$. Then the gap sequence is [2 1 0 0 1][2\ 1\ 0\ 0\ 1]. It follows that $t_{1}^{3}\oplus 0\in R$, $t_{1}^{5}\oplus t_{2}^{3}\in R$ (up to rescaling $t_{1}$), $t_{1}^{4}\oplus 0\in R$, and $0\oplus t_{2}^{4}\in R$. Since $t_{1}^{5}\oplus 0\not\in R$, the quadratic generator of $\mathfrak{m}/\mathfrak{m}^{2}$ must be $0\oplus t_{2}^{2}$.
@@
    In terms of these generators, the defining equations of this singularity are (xy, xw, yz - w^2, zw - y^3, x^3 + y^2w - z^2).
    Since $x$, $y$, $z$, and $w$ have weights $10$, 99, $15$, and $12$ under the $\mathbb{G}_{m}$-action, respectively, the standard projectivization of the corresponding Gorenstein curve $Y$ as well as its deformations $(C,p_{1},p_{2})$ can be embedded in the weighted projective space $\mathbb{P}(10,9,15,12,1)$, where the divisor at infinity is cut out by the hyperplane of weight $1$ as $\frac{1}{3}p_{1}+\frac{1}{5}p_{2}$.
@@
    Finally, we compute the weights and characters for this singularity. In this case, we have $\ell=15$, $a_{1}=3$, and $a_{2}=5$. Since $\omega_{C}^{\log}\sim 5p_{1}+3p_{2}$, it follows that F^λ H^0(C, ω_C^log) = H^0(C, (5 - ⌈λ/3⌉)p_1 + (3 - ⌈λ/ 5⌉) p_2). Therefore, the filtration of $H^{0}(C,\omega_{C}^{\log})$ is given by
@@
    The nontrivial weights are $3$, 55, 66, and $15$. We thus obtain that χ_1^log(4,2)^odd = 29,  χ_2^log(4,2)^odd = 149,  α(4,2)^odd = 79228.
@@ 5.6. The stratum $\mathbb{P}\mathcal{H}(3,3)^{\operatorname{nonhyp}}$
In this case, we have $g=4$, $b=2$, and $\delta=5$. The gap sequence is determined by $\alpha_{1}\ \alpha_{2}\ \alpha_{3}\ 1]$, where $\alpha_{4}=1$ is the last nonzero entry. Moreover, $\sum_{i\geq 1}\alpha_{i}=4$, $\alpha_{1}\neq 0$, and $\alpha_{2}\neq 0$. Therefore, the gap sequence has only two possibilities: [1 1 1 1][1\ 1\ 1\ 1] or [2 1 0 1][2\ 1\ 0\ 1]. Note that $\alpha_{1}=1$ or $2$ corresponds to $\dim H^{0}(Y,y_{1}+y_{2})=2$ or $1$, respectively. Since here we consider the nonhyperelliptic component, its gap sequence is [2 1 0 1][2\ 1\ 0\ 1].
@@
This singularity is denoted by $Z_{9}$ in Giusti’s classification of simple space curve singularities; see [Loo84a, (7.22)] and [Ste15, Table 2].
Since $x$, $y$, and $z$ have weights $2$, 33, and $3$ under the $\mathbb{G}_{m}$-action, respectively, the standard projectivization of the corresponding Gorenstein curve $Y$ as well as its deformations $(C,p_{1},p_{2})\in\mathbb{P}\mathcal{H}(3,3)^{\operatorname{nonhyp}}$ can be embedded in the weighted projective space $\mathbb{P}(2,3,3,1)$, where the divisor at infinity is cut out by the hyperplane of weight $1$ as $p_{1}+p_{2}$.
@@
The nontrivial weights are $1$, 11, 22, and $4$. We thus obtain that
@@ 5.7. The stratum $\mathbb{P}\mathcal{H}(3,2,1)$
In this case, we have $g=4$, $b=3$, and $\delta=6$. The gap sequence is determined by $\alpha_{1}\ \alpha_{2}\ \alpha_{3}\ 1]$, where $\alpha_{4}=1$ is the last nonzero entry. Moreover, $\sum_{i\geq 1}\alpha_{i}=4$ and $\alpha_{1}\geq\alpha_{2}>0$. Therefore, the gap sequence is [2 1 0 1][2\ 1\ 0\ 1].
@@
Since $x$, $y$, $z$, and $w$ have weights $6$, 88, 99, and $12$ under the $\mathbb{G}_{m}$-action, respectively, the standard projectivization of the corresponding Gorenstein curve $Y$ as well as its deformations $(C,p_{1},p_{2},p_{3})\in\mathbb{P}\mathcal{H}(3,2,1)$ can be embedded in the weighted projective space $\mathbb{P}(6,8,9,12,1)$, where the divisor at infinity is cut out by the hyperplane of weight $1$ as $\frac{1}{3}p_{1}+\frac{1}{4}p_{2}+\frac{1}{6}p_{3}$.
@@
The nontrivial weights are $3$, 44, 66, and $12$. We thus obtain that
@@ 5.8. The stratum $\mathbb{P}\mathcal{H}(2,2,2)^{\operatorname{odd}}$
In this case, we have $g=4$, $b=3$, and $\delta=6$. The gap sequence is determined by $\alpha_{1}\ \alpha_{2}\ 1]$, where $\alpha_{3}=1$ is the last nonzero entry. Moreover, $\sum_{i\geq 1}\alpha_{i}=4$, $\alpha_{1}\neq 0$, and $\alpha_{1}\geq\alpha_{2}$. The gap sequence has two possibilities: [2 1 1][2\ 1\ 1] or [3 0 1][3\ 0\ 1]. Note that $\alpha_{1}=2$ or $3$ corresponds to $\dim H^{0}(Y,y_{1}+y_{2}+y_{3})=2$ or $1$, respectively. Since here we consider the odd spin component, its gap sequence is [3 0 1][3\ 0\ 1].
@@
Since $x$, $y$, $z$, $w$, and $u$ have weights $2$, 22, 22, 33, and $3$ under the $\mathbb{G}_{m}$-action, respectively, the standard projectivization of the corresponding Gorenstein curve $Y$ as well as its deformations $(C,p_{1},p_{2},p_{3})\in\mathbb{P}\mathcal{H}(2,2,2)^{\operatorname{odd}}$ can be embedded in the weighted projective space $\mathbb{P}(2,2,2,3,3,1)$, where the divisor at infinity is cut out by the hyperplane of weight $1$ as $p_{1}+p_{2}+p_{3}$.
@@
The nontrivial weights are $1$, 11, 11, and $3$. We thus obtain that